If the force on a rocket moving with a velocity of 300 m/sec is 345N, then the rate of combustion of the fuel, is


        (a)    0.55 kg/sec


        (b)    0.75 kg/sec


        (c)    1.15 kg/sec       


(d)    2.25 kg/sec


Grade:11

2 Answers

Navjyot Kalra
10 years ago

(c)
Velocity of the rocket (u) = 300 m/s and force (F) = 345N.
Rate of combustion of fuel
    (dm/dt)=F/u=1.15 kg/sec    

Hrishant Goswami
10 years ago

(c)

Velocity of the rocket (u) = 300 m/s and force (F) = 345N.

Rate of combustion of fuel